OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 3y = 3 3x – 2y = 11 What is the solution to the system? (–3, 3) (–1, –7) (1, –4) (3, –1)

hero (hero):

y = (3 - 2x)/3 y = (3x-11)/2 graph using graphing calc

OpenStudy (anonymous):

lol graphing technology.

OpenStudy (anonymous):

HOW DO I DO IT ? IM NEW TO THIS

OpenStudy (anonymous):

Okay. First i assume you know how to use a graphing calculator. the GC only accpets equation that are in the form of y=Ax.... and so on. The impt part is the "y=" So as what Hero say, you express the equation in that form to get, y = (3 - 2x)/3 y = (3x-11)/2 Then put these 2 eqn into the GC, Then find the intercept of the 2 lines. that will give u the solution

OpenStudy (anonymous):

I STILL DONT UNDERSTAND , BUT THANKS ANYWAY